OpenStudy (anonymous):

Graphs are not my strong suit, is anyone on here able to help me? :) I'll answer yours in return!!!! Choose the system of equations with the graphical solution shown below: 5x - y = 612x + 6y = 6 5x + y = -612x + 6y = 6 5x - y = 612x - 6y = 6 5x + y = -612x - 6y = -6

OpenStudy (anonymous):

This is the graph:

OpenStudy (anonymous):

the graphs meets at (1,-1) right ?

OpenStudy (anonymous):

Yes sir

OpenStudy (anonymous):

so lets see which one of the sets satisfies (1,-1)

OpenStudy (anonymous):

Either B or D

OpenStudy (anonymous):

A satisfies 5x - y = 6 12x + 6y = 6 B not : 5x + y = -6 this is not correct : 5 - (-1) = -6 12x + 6y = 6 D not: for the same reason 5x + y = -6 12x - 6y = -6

OpenStudy (anonymous):

for A : 5 * 1 - (-1) = 6 correct 12 * 1 + 6 * (-1) = 6 correct

OpenStudy (anonymous):

You are VERY good at this....

OpenStudy (anonymous):

you just have to plug in the values carefully and do the math slowly so you want do mistakes with the signs

OpenStudy (anonymous):

Do you mind if I write this problem down so i can learn it and refer back to it?

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

I have a tendency of being thrown off by signs

OpenStudy (anonymous):

Thank you so much @Coolsector

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

so you do it here ?